PLEASE SHOW ALL OF YOUR WORK CLEARLY, STEP-BY-STEP!

1. The manager of Moxa, a local coffee shop, is interested in bringing in more customers so he hired an YU student to hand out fliers on campus to try to reach the student body. The owner of the shop wanted to know how much it would cost to bring each new customer into the shop before she signed off on the plan. Using the information below, please calculate the customer acquisition cost for Moxa (the cost to acquire each customer)

Cost for fliers= $.25 / unit

Cost for student to hand out fliers = $1000 for entire campaign

Number of students who will be given a flier = 25,000

Based on information provided by the manager of Subway, the restaurant next door, the Moxa manager determined that 80% of the students who received a flier throw it in the trash. Of those who actually kept it, 1/4 read it. And of those students who kept and read the flier, 10% actually stopped by the shop and purchased a beverage.
Answer #1

Total Cost of Acquisition= Cost of fliers*No. of Students + Cost for student to hand out fliers

= 0.25*25000+1000

= $ 7250

Calculating the no of Student who will buy from shop:

No. of Student who will not throw flier in trash = 25000*(1-80%)

= 5000 students

No. of students who kept it & read it = 5000*1/4

= 1250

No of Student who will buy from shop = 1250*10%

= 125 students

So, the cost to acquire each customer = Total Cost/No. of students who will buy from shop

= 7250/125

= $ 58

So, The cost to acquire per Customer is $ 58

Note- If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.